# ChemMLLM: Chemical Multimodal Large Language Model

### v1

@@ Appendix F Evaluation Metrics
(1) img2caption: We use BLEU-2/4, ROUGE-1/2/L, and METEOR to evaluate the quality of generated captions against reference texts; (2) img2property: We extract seven molecular properties from LLM-generated outputs and evaluate the accuracy using Mean Squared Error (MSE), Mean Absolute Error (MAE), and Pearson correlation; (3) img2smiles: We extract SMILES strings from the LLM outputs and adopt Tanimoto similarity and Tanimoto hit 1.0 (tanimoto@1.0) that measures the percentage of exact matches (similarity = 1.0); (4) property2img: Generated molecular images are converted to SMILES via MolScribe Qian et al. (2023), from which properties are computed and evaluated using MSE, MAE, and Pearson correlation. Each model is run five times, and the best result is reported; (5) img2img: We use Increased LogP, as well as molecular diversity and novelty to measure the optimized molecule. Other settings are the same as property2img.
@@
    where $p_{n}$ denotes the modified n-gram precision for n-grams of size $n$, and BP is the brevity penalty, which penalizes short candidate sentences to prevent artificially high scores. The BLEU-N score ranges from 0 to 1, where a higher score indicates better overlap with the reference text in terms of n-gram content. A higher BLEU-N value generally reflects better fluency and adequacy in the generated text. In our experimental evaluation, we use the word_tokenize() function from the NLTK library  Bird et al. (2009) to do tokenization and employ the sentence_bleu() metric with uniform weights for all n-gram precision calculations (i.e., equal weights for 1- to 4-gram contributions).
@@
    where $n$ denotes the length of the n-grams (e.g., ROUGE-1 for unigrams, ROUGE-2 for bigrams), and $\text{Count}_{\text{match}}(\text{gram}_{n})$ is the number of n-grams in the reference that also appear in the candidate text. ROUGE-N values range from 0 to 1 and a higher ROUGE-N value indicates better performance. In our experimental evaluation, we employ the rouge_scorer() metric from the rouge_score Lin (2004) library.
@@
    where $F_{\text{mean}}=\frac{10\cdot P\cdot R}{R+9P}$, with $P$ and $R$ denoting unigram precision and recall, respectively. The penalty is a function of the number of chunks in the alignment, designed to penalize disordered matches. METEOR scores range from 0 to 1, with higher scores indicating better alignment with the reference text. In our experimental evaluation, we perform tokenization using the word_tokenize() function and employ the METEOR metric (meteor_score()), both implemented in the NLTK library Bird et al. (2009).
@@
    $\text{diversity}(\mathcal{Z})=$ $1-\frac{1}{|\mathcal{Z}|(|\mathcal{Z}|-1)}\cdot$ (15)
    $\sum_{X,Y\in\mathcal{Z},\newline X\neq Y}\text{sim}(X,Y),$
@@
  • Tanimoto@1 is to measures the proportion of generated molecules that exactly match the ground-truth molecule in terms of Tanimoto similarity. Specifically, it computes the ratio of generated molecules whose Tanimoto similarity with the corresponding ground-truth molecule is equal to 1. It is defined as
[…]
    $\mathbb{I}\left[\text{Tanimoto}(f_{i}^{\text{gen}},f_{i}^{\text{true}})=1\right],$
    where $N$ is the number of generated molecules, $f_{i}^{\text{gen}}$ and $f_{i}^{\text{true}}$ are the Morgan fingerprints of the $i$-th generated and ground-truth molecules, respectively, and $\mathbb{I}[\cdot]$ is the indicator function. The score ranges from 0 to 1, where a higher Tanimoto@1 indicates better exact matching performance between generated and reference molecules.
@@
    Increased $\text{Increased LogP}=\frac{1}{N}\sum_{i=1}^{N}$ (17)
    $\left(\text{LogP}(m_{i}^{\text{opt}})-\text{LogP}(m_{i}^{\text{orig}})\right),$
@@
    $\left(1-\text{Tanimoto}(f_{i},f_{j})\right),$
    where $N$ is the number of molecules in the set, and $f_{i}$ and $f_{j}$ are the Morgan fingerprints of the $i$-th and $j$-th molecules, respectively. The Tanimoto similarity measures the overlap between two binary fingerprints, and the distance is computed as $1-\text{Tanimoto}$. The diversity values range from 0 to 1, with higher values indicating greater chemical diversity.
@@
    Valid $\text{Valid Rate}=\frac{1}{N}\sum_{i=1}^{N}$ (20)
    $\mathbb{I}\left[\text{structured}(o_{i})\land\text{valid}(o_{i})\right],$
    where $N$ is the total number of model outputs, $o_{i}$ is the $i$-th output, $\text{structured}(\cdot)$ checks whether the output follows the expected structured format, and $\text{valid}(\cdot)$ verifies the syntactic validity of any SMILES strings present in the output. The score ranges from 0 to 1, with a higher Validindicates better adherence to the required output format and chemical validity.
@@ Appendix G Molecular Properties
  • TPSA: Topological Polar Surface Area (TPSA) quantifies the surface area contributed by polar atoms, typically oxygen and nitrogen, including their attached hydrogens. The theoretical TPSA ranges from 00 to several hundreds or even thousands of Å2 for highly polar or large biomolecules.
@@ Appendix H Implementation Details
For the first training stage, Mol-VQGAN is trained on 8 NVIDIA A800×80G GPUs for two epochs. The batch size is set to 16 and the base learning rate is set to 4.5e-06. We use Adam Kingma and Ba (2014) as optimizer and $\mathcal{L}_{rec}+\lambda_{1}\mathcal{L}_{perceptual}$ as monitor to validate the model and save the best checkpoint during 2-epoch training. For the second training stage, ChemMLLM is trained on 8 NVIDIA A800×80G GPUs for three epochs. The batch size is set to be 16, the base learning rate is set to be 2e-5 and z-loss weight is set to be 1e-5. We use AdamW Loshchilov and Hutter (2017) as optimizer and employ mixed-precision training, utilizing brain floating point 16 precision (bf16) Kalamkar et al. (2019) for forward propagation and 32-bit floating point precision (fp32) for backward propagation to balance the training efficiency and stability. To handle distributed training, we apply PyTorch Fully Shared Data Parallel (FSDP) Zhao et al. (2023) strategy. We train three variants of ChemMLLM for different tasks.
@@
For Mol-VQGAN training code, we use the official implementation code of original VQGAN Esser et al. (2021). Specifically, we utilize the synthesized image datasets to let the original Chameleon VQGAN learn how to understand and generate molecule images. We sample 1,000,000 molecule images from PubChem synthesized by RDKit Landrum et al. (2006) and combine them with all images synthesized in 5 downstream tasks as training dataset for Mol-VQGAN. All parameters of Encoder, Decoder and codebook of VQGAN are trained. We first train VQGAN on this 1 million-level molecule image dataset for two epochs and save the best check point according to $\mathcal{L}_{rec}+\lambda_{1}\mathcal{L}_{perceptual}$. After the initial training, we find that it can not reconstruct image well on the dataset with less data size like img2caption dataset, so we do continuous training based on the best checkpoint using small-size datasets. Specifically, we continue training the 1-million best check point on img2caption images for five epochs and save the best checkpoint. Finally, we get the well-trained Mol-VQGAN to tokenize molecule images.The original VQGAN will result in unclear and distorted images when encoding and decoding molecule images. After training, Mol-VQGAN can encode and decode molecule image almost the same with original image. Several examples are shown in Figure 15. The validation curve from start to the best checkpoint is shown in Figure 14.
@@
Figure 15: Examples of origin VQGAN and Mol-VQGAN. The origin images is shown in the first row. As shown in the second row, the origin VQGAN can not encode and decode molecule image clearly and accurately, resulting in distorted atoms and bonds which are hard to distinguish. As shown in the third row, our well-trained Mol-VQGAN can encode and decode molecule images with clear atoms and bonds, which is almost the same as the origin molecule images.
@@
ChemMLLM Supervised FineTuning Training For ChemMLLM training code, we utilize Lumina-mGPT framework Liu et al. (2024). We train three variants ChemMLLM  ChemMLLM-pro2img and ChemMLLM-img2img.
@@
For ChemMLLM, we train it on img2caption, img2property, img2smiles and property2img datasets. Apart from the four datasets, we also train it on SMILES molecule image generation (smiles2img) and text based drug design dataset (caption2img), which is the inverse task of img2smiles and img2caption. smiles2img requires model to receive a SMILES and output the corresponding image and caption2img ask model to output molecule image according to molecule caption. We train ChemMLLM on the six datasets for three epochs; For property2img task, we train and evaluate ChemMLLM-pro2img. We find that directly training the model on raw data can not achieve good performance, so we do data augmentation by rotating images by $90^{\circ}$,$180^{\circ}$, and $270^{\circ}$ so as to generate augmented data that is four times the size of the original dataset. We train ChemMLLM-pro2img on the augmented dataset for two epochs; For img2img task, we train and evaluate ChemMLLM-img2img. we train it on img2img dataset for two epochs. The training hyper-parameters primarily follow Lumina-mGPT Liu et al. (2024).
@@
Model task # Training # Test
ChemMLLM img2caption 69,799 3,300
img2property 95,000 5,000
img2smiles 95,000 5,000
property2img 95,000 5,000
smiles2img 95,000 5,000
caption2img 72,143 3301
ChemMLLM-pro2img property2img 380,000 20,000
ChemMLLM-img2img img2img 157,673 17,520
@@
Settings ChemMLLM ChemMLLM-pro2img ChemMLLM-img2img
z-loss weight 1e-05 1e-05 1e-05
warmup epochs 0.01 0.01 0.01
learning rate 2e-05 2e-05 2e-05
weight decay 0.1 0.1 0.1
drop rate 0.05 0.05 0.05
total bacth size $16\times 8\times 1$ $16\times 8\times 1$ $8\times 4\times 1$
GPUs for training 8×8$\times$8 ×A800 (80G) 8×8$\times$8 × A800 (80G) 4×4$\times$4 × A800 (80G)
GPUs hours(h) 65 30.6 25.4

### v2

@@ Appendix F Evaluation Metrics
MW LogP TPSA
Chemception (CNN) 0.87±0.01$\pm$ 0.01± 0.01 4180.95±446.65$\pm$ 446.65± 446.65 44.75±1.48$\pm$ 1.48± 1.48 0.62 ±0.02$\pm$ 0.02± 0.02 2.95±0.34$\pm$ 0.34± 0.34 1.23 ±0.03$\pm$ 0.03± 0.03 0.84±0.03$\pm$ 0.03± 0.03 768.45±297.01$\pm$ 297.01± 297.01 14.45±0.74$\pm$ 0.74± 0.74 100%
Chemprop (GNN) 0.99 ±0.0003$\pm$ 0.0003± 0.0003 21.32±10.70$\pm$ 10.70± 10.70 1.63 ±0.13$\pm$ 0.13± 0.13 0.98 ±0.003$\pm$ 0.003± 0.003 0.10 ±0.03$\pm$ 0.03± 0.03 0.16±0.009$\pm$ 0.009± 0.009 0.99 ±0.0001$\pm$ 0.0001± 0.0001 4.40 ±1.69$\pm$ 1.69± 1.69 0.91 ±0.05$\pm$ 0.05± 0.05 100%
Qwen-VL-Chat (7B) 0.78 ±0.02$\pm$ 0.02± 0.02 7073.35 ±856.93$\pm$ 856.93± 856.93 58.07 ±1.94$\pm$ 1.94± 1.94 0.14 ±0.02$\pm$ 0.02± 0.02 4.43 ±0.45$\pm$ 0.45± 0.45 1.48 ±0.04$\pm$ 0.04± 0.04 0.19 ±0.02$\pm$ 0.02± 0.02 $>$1e+4 ±547.06$\pm$ 547.06± 547.06 82.83 ±1.88$\pm$ 1.88± 1.88 99.3%
InternVL-Chat-v1.5 (20B) 0.59 ±0.02$\pm$ 0.02± 0.02 $>$1e+4 ±1588.37$\pm$ 1588.37± 1588.37 83.60±3.63$\pm$ 3.63± 3.63 0.04±0.05$\pm$ 0.05± 0.05 9.24±0.62$\pm$ 0.62± 0.62 2.37±0.08$\pm$ 0.08± 0.08 0.29±0.03$\pm$ 0.03± 0.03 2158.86±433.77$\pm$ 433.77± 433.77 28.73 ±1.52$\pm$ 1.52± 1.52 55.0%
[…]
ChemVLM-8B 0.84 ±0.02$\pm$ 0.02± 0.02 9573.42 ±1961.79$\pm$ 1961.79± 1961.79 56.94 ±4.46$\pm$ 4.46± 4.46 0.38 ±0.05$\pm$ 0.05± 0.05 4.90 ±0.58$\pm$ 0.58± 0.58 1.68 ±0.08$\pm$ 0.08± 0.08 0.26 ±0.06$\pm$ 0.06± 0.06 5332.09 ±757.56$\pm$ 757.56± 757.56 53.66 ±2.78$\pm$ 2.78± 2.78 31.3%
ChemMLLM-7B (ours) 0.97±0.004$\pm$ 0.004± 0.004* 789.72±162.85$\pm$ 162.85± 162.85* 16.17±0.72$\pm$ 0.72± 0.72* 0.92±0.01$\pm$ 0.01± 0.01* 0.70±0.14$\pm$ 0.14± 0.14* 0.52±0.02$\pm$ 0.02± 0.02* 0.97±0.005$\pm$ 0.005± 0.005* 152.65±39.26$\pm$ 39.26± 39.26* 6.06±0.33$\pm$ 0.33± 0.33* 99.6%
ChemMLLM-34B (ours) 0.98±0.002$\pm$ 0.002± 0.002* 419.95±86.08$\pm$ 86.08± 86.08* 11.57±0.53$\pm$ 0.53± 0.53* 0.93±0.01$\pm$ 0.01± 0.01* 0.53±0.11$\pm$ 0.11± 0.11* 0.43±0.01$\pm$ 0.01± 0.01* 0.98±0.004$\pm$ 0.004± 0.004* 88.49±26.60$\pm$ 26.60± 26.60* 3.54±0.28$\pm$ 0.28± 0.28* 99.8%
Hbd Hba Rb QED
Chemception(CNN) 0.81±0.02$\pm$ 0.02± 0.02 1.03±0.22$\pm$ 0.22± 0.22 0.58±0.02$\pm$ 0.02± 0.02 0.84±0.01$\pm$ 0.01± 0.01 1.92±0.29$\pm$ 0.29± 0.29 0.93±0.03$\pm$ 0.03± 0.03 0.73±0.02$\pm$ 0.02± 0.02 7.57±0.79$\pm$ 0.79± 0.79 1.91±0.06$\pm$ 0.06± 0.06 -0.004±0.03$\pm$ 0.03± 0.03 0.25±0.01$\pm$ 0.01± 0.01 0.40±0.009$\pm$ 0.009± 0.009
Chemprop(GNN) 0.99 ±8.68$\pm$ 8.68± 8.68 0.03 ±0.0009$\pm$ 0.0009± 0.0009 0.03 ±0.001$\pm$ 0.001± 0.001 0.99 ±0.0004$\pm$ 0.0004± 0.0004 0.03 ±0.005$\pm$ 0.005± 0.005 0.11 ±0.004$\pm$ 0.004± 0.004 0.99 ±0.0002$\pm$ 0.0002± 0.0002 0.05 ±0.005$\pm$ 0.005± 0.005 0.15 ±0.005$\pm$ 0.005± 0.005 0.89 ±0.006$\pm$ 0.006± 0.006 0.009 ±0.0005$\pm$ 0.0005± 0.0005 0.07 ±0.001$\pm$ 0.001± 0.001
Qwen-VL-Chat (7B) -0.02 ±0.008$\pm$ 0.008± 0.008 4.87 ±0.75$\pm$ 0.75± 0.75 1.42±0.05$\pm$ 0.05± 0.05 0.05±0.01$\pm$ 0.01± 0.01 30.56±1.24$\pm$ 1.24± 1.24 4.91±0.08$\pm$ 0.08± 0.08 0.19±0.10$\pm$ 0.10± 0.10 145.40±31.19$\pm$ 31.19± 31.19 6.87±0.31$\pm$ 0.31± 0.31 0±0.0$\pm$ 0.0± 0.0 - -
InternVL-Chat-v1.5 (20B) 0.03±0.05$\pm$ 0.05± 0.05 9.92±0.89$\pm$ 0.89± 0.89 2.24±0.09$\pm$ 0.09± 0.09 0.22±0.04$\pm$ 0.04± 0.04 10.24±0.85$\pm$ 0.85± 0.85 2.41±0.08$\pm$ 0.08± 0.08 $\pm 0.04$ 43.67±4.28$\pm$ 4.28± 4.28 4.82±0.19$\pm$ 0.19± 0.19 0.$\pm 0.03$ 4.01±3.26$\pm$ 3.26± 3.26 0.40±0.08$\pm$ 0.08± 0.08
LLaVA-v1.5-7B 0.004 ±0.05$\pm$ 0.05± 0.05 53.31±27.73$\pm$ 27.73± 27.73 3.74±0.33$\pm$ 0.33± 0.33 0.04±0.05$\pm$ 0.05± 0.05 23.71±5.44$\pm$ 5.44± 5.44 2.93±0.19$\pm$ 0.19± 0.19 0.03 ±0.04$\pm$ 0.04± 0.04 39.99±10.63$\pm$ 10.63± 10.63 3.85±0.26$\pm$ 0.26± 0.26 -0.11±0.08$\pm$ 0.08± 0.08 >>$\pm>1e+5$ 24.53±18.53$\pm$ 18.53± 18.53
ChemVLM-8B 0.49 ±0.09$\pm$ 0.09± 0.09 4.25 ±0.85$\pm$ 0.85± 0.85 1.35 ±0.08$\pm$ 0.08± 0.08 0.32 ±0.07$\pm$ 0.07± 0.07 27.27 ±1.95$\pm$ 1.95± 1.95 4.58 ±0.14$\pm$ 0.14± 0.14 0.10 ±0.06$\pm$ 0.06± 0.06 45.88 ±5.02$\pm$ 5.02± 5.02 5.56 ±0.22$\pm$ 0.22± 0.22 -0.003 ±0.06$\pm$ 0.06± 0.06 0.24 ±0.02$\pm$ 0.02± 0.02 0.37±0.01$\pm$ 0.01± 0.01
ChemMLLM-7B (ours) 0.96±0.004$\pm$ 0.004± 0.004* 0.18±0.02$\pm$ 0.02± 0.02* 0.13±0.01$\pm$ 0.01± 0.01* 0.94±0.007$\pm$ 0.007± 0.007* 0.79±0.12$\pm$ 0.12± 0.12* 0.44±0.02$\pm$ 0.02± 0.02* 0.94±0.01$\pm$ 0.01± 0.01* 1.62±0.33$\pm$ 0.33± 0.33* 0.59±0.03$\pm$ 0.03± 0.03* 0.91±0.006$\pm$ 0.006± 0.006* 0.008±0.0004$\pm$ 0.0004± 0.0004* 0.06±0.002$\pm$ 0.002± 0.002*
ChemMLLM-34B (ours) 0.96±0.01$\pm$ 0.01± 0.01* 0.22±0.06$\pm$ 0.06± 0.06* 0.11±0.01$\pm$ 0.01± 0.01* 0.96±0.003$\pm$ 0.003± 0.003* 0.53±0.12$\pm$ 0.12± 0.12* 0.26±0.02$\pm$ 0.02± 0.02* 0.97±0.006$\pm$ 0.006± 0.006* 0.75±0.20$\pm$ 0.20± 0.20* 0.37±0.02$\pm$ 0.02± 0.02* 0.93±0.005$\pm$ 0.005± 0.005* 0.006±0.0005$\pm$ 0.0005± 0.0005* 0.05±0.001$\pm$ 0.001± 0.001*
Table 15: Results on img2property task: Hbd, Hba, Rb, and QED (best, 2nd best, *: significantly better). Task-specific models Chemception (Goh et al. 2017) and Chemprop (Heid et al. 2023) take molecular image and graph as features, respectively.
Qwen-VL-Chat (7B) 0.09 ±0.001$\pm$ 0.001± 0.001 0.01 ±0.0009$\pm$ 0.0009± 0.0009 0.32 ±0.003$\pm$ 0.003± 0.003 0.07 ±0.001$\pm$ 0.001± 0.001 0.22 ±0.002$\pm$ 0.002± 0.002 0.19 ±0.001$\pm$ 0.001± 0.001
InternVL-Chat-v1.5 (20B) 0.04±0.0008$\pm$ 0.0008± 0.0008 0.001 ±0.0002$\pm$ 0.0002± 0.0002 0.38±0.003$\pm$ 0.003± 0.003 0.07 ±0.001$\pm$ 0.001± 0.001 0.26 ±0.002$\pm$ 0.002± 0.002 0.19±0.001$\pm$ 0.001± 0.001
LLaVA-v1.5-7B 0.08 ±0.001$\pm$ 0.001± 0.001 0.004 ±0.0004$\pm$ 0.0004± 0.0004 0.33 ±0.002$\pm$ 0.002± 0.002 0.07 ±0.001$\pm$ 0.001± 0.001 0.23 ±0.002$\pm$ 0.002± 0.002 0.20 ±0.001$\pm$ 0.001± 0.001
GPT-4o 0.16 ±0.003$\pm$ 0.003± 0.003 0.07±0.002$\pm$ 0.002± 0.002 0.28±0.005$\pm$ 0.005± 0.005 0.13±0.003$\pm$ 0.003± 0.003 0.23±0.004$\pm$ 0.004± 0.004 0.22±0.004$\pm$ 0.004± 0.004
ChemVLM-8B 0.15 ±0.004$\pm$ 0.004± 0.004 0.08 ±0.003$\pm$ 0.003± 0.003 0.28±0.006$\pm$ 0.006± 0.006 0.13 ±0.003$\pm$ 0.003± 0.003 0.23 ±0.005$\pm$ 0.005± 0.005 0.23 ±0.005$\pm$ 0.005± 0.005
ChemMLLM-7B (ours) 0.33 ±0.005$\pm$ 0.005± 0.005* 0.21±0.005$\pm$ 0.005± 0.005* 0.50±0.005$\pm$ 0.005± 0.005* 0.31±0.006$\pm$ 0.006± 0.006* 0.43±0.005$\pm$ 0.005± 0.005* 0.43±0.005$\pm$ 0.005± 0.005*
ChemMLLM-34B (ours) 0.36 ±0.006$\pm$ 0.006± 0.006* 0.25±0.006$\pm$ 0.006± 0.006* 0.51±0.006$\pm$ 0.006± 0.006* 0.33±0.006$\pm$ 0.006± 0.006* 0.45±0.006$\pm$ 0.006± 0.006* 0.45±0.006$\pm$ 0.006± 0.006*
@@
(1) img2caption: We use BLEU-2/4, ROUGE-1/2/L, and METEOR to evaluate the quality of generated captions against reference texts; (2) img2property: We extract seven molecular properties from LLM-generated outputs and evaluate the accuracy using Mean Squared Error (MSE), Mean Absolute Error (MAE), and Pearson correlation; (3) img2smiles: We extract SMILES strings from the LLM outputs. Then, we use accuracy that measures exact matching (based on canonical SMILES) and Tanimoto similarity that measures similarities between generated and groundtruth SMILES; (4) property2img: Generated molecular images are converted to SMILES via MolScribe (Qian et al. 2023), from which properties are computed and evaluated using MSE, MAE, and Pearson correlation. Each model is run five times, and the best result is reported; (5) img2img: We use Increased LogP, as well as molecular diversity and novelty to measure the optimized molecule. Other settings are the same as property2img.
@@
    where $p_{n}$ denotes the modified n-gram precision for n-grams of size $n$, and BP is the brevity penalty, which penalizes short candidate sentences to prevent artificially high scores. The BLEU-N score ranges from 0 to 1, where a higher score indicates better overlap with the reference text in terms of n-gram content. A higher BLEU-N value generally reflects better fluency and adequacy in the generated text. In our experimental evaluation, we use the word_tokenize() function from the NLTK library (Bird, Klein, and Loper 2009) to do tokenization and employ the sentence_bleu() metric with uniform weights for all n-gram precision calculations (i.e., equal weights for 1- to 4-gram contributions).
@@
    where $n$ denotes the length of the n-grams (e.g., ROUGE-1 for unigrams, ROUGE-2 for bigrams), and $\text{Count}_{\text{match}}(\text{gram}_{n})$ is the number of n-grams in the reference that also appear in the candidate text. ROUGE-N values range from 0 to 1 and a higher ROUGE-N value indicates better performance. In our experimental evaluation, we employ the rouge_scorer() metric from the rouge_score (Lin 2004) library.
@@
    where $F_{\text{mean}}=\frac{10\cdot P\cdot R}{R+9P}$, with $P$ and $R$ denoting unigram precision and recall, respectively. The penalty is a function of the number of chunks in the alignment, designed to penalize disordered matches. METEOR scores range from 0 to 1, with higher scores indicating better alignment with the reference text. In our experimental evaluation, we perform tokenization using the word_tokenize() function and employ the METEOR metric (meteor_score()), both implemented in the NLTK library (Bird, Klein, and Loper 2009).
@@
    $\text{diversity}(\mathcal{Z})$ (15)
    $=$ $1-\frac{1}{|\mathcal{Z}|(|\mathcal{Z}|-1)}\cdot\sum_{X,Y\in\mathcal{Z},\newline X\neq Y}\text{sim}(X,Y),$
@@
  • Accuracy refers to the proportion of generated molecules that exactly match the ground-truth molecule. Specifically, it computes the ratio of generated molecules whose canonical SMILES is the same as the canonical SMILES of the ground-truth. It is defined as
    $\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left[\text{Match}(s_{i}^{\text{gen}},s_{i}^{\text{true}})\right],$ (16)
    where $N$ is the number of generated molecules, $s_{i}^{\text{gen}}$ and $s_{i}^{\text{true}}$ are the canonical smiles of the $i$-th generated and ground-truth molecules, respectively, and $\mathbb{I}[\cdot]$ is the indicator function. The score ranges from 0 to 1, where a higher Accuracy indicates better exact matching performance between generated and reference molecules.
@@
    Increased LogP (17)
    $=$ $\frac{1}{N}\sum_{i=1}^{N}\left(\text{LogP}(m_{i}^{\text{opt}})-\text{LogP}(m_{i}^{\text{orig}})\right),$
@@
    Diversity (18)
    $=$ $\frac{2}{N(N-1)}\sum_{i=1}^{N}\sum_{j=i+1}^{N}\left(1-\text{Tanimoto}(f_{i},f_{j})\right),$
@@
    $\text{Valid Rate}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left[\text{structured}(o_{i})\land\text{valid}(o_{i})\right],$ (20)
[…]
@@ Appendix G Molecular Properties
  • TPSA: Topological Polar Surface Area (TPSA) quantifies the surface area contributed by polar atoms, typically oxygen and nitrogen, including their attached hydrogens. The theoretical TPSA ranges from 0 to several hundreds or even thousands of Å2 for highly polar or large biomolecules.
@@ Appendix H Implementation Details
For the first training stage, Mol-VQGAN is trained on 8 NVIDIA A800×80G GPUs for two epochs. The batch size is set to 16 and the base learning rate is set to 4.5e-06. We use Adam (Kingma and Ba 2014) as optimizer and $\mathcal{L}_{rec}+\lambda_{1}\mathcal{L}_{perceptual}$ as monitor to validate the model and save the best checkpoint during 2-epoch training. For the second training stage, ChemMLLM is trained on 8 NVIDIA A800×80G GPUs for three epochs. The batch size is set to be 16, the base learning rate is set to be 2e-5 and z-loss weight is set to be 1e-5. We use AdamW (Loshchilov and Hutter 2017) as optimizer and employ mixed-precision training, utilizing brain floating point 16 precision (bf16) (Kalamkar et al. 2019) for forward propagation and 32-bit floating point precision (fp32) for backward propagation to balance the training efficiency and stability. To handle distributed training, we apply PyTorch Fully Shared Data Parallel (FSDP) (Zhao et al. 2023) strategy. We train three variants of ChemMLLM for different tasks.
@@
For Mol-VQGAN training code, we use the official implementation code of original VQGAN (Esser, Rombach, and Ommer 2021). Specifically, we utilize the synthesized image datasets to let the original Chameleon VQGAN learn how to understand and generate molecule images. We sample 1,000,000 molecule images from PubChem synthesized by RDKit (Landrum et al. 2006) and combine them with all images synthesized in 5 downstream tasks as training dataset for Mol-VQGAN. All parameters of Encoder, Decoder and codebook of VQGAN are trained. We first train VQGAN on this 1 million-level molecule image dataset for two epochs and save the best check point according to $\mathcal{L}_{rec}+\lambda_{1}\mathcal{L}_{perceptual}$. After the initial training, we find that it can not reconstruct image well on the dataset with less data size like img2caption dataset, so we do continuous training based on the best checkpoint using small-size datasets. Specifically, we continue training the 1-million best check point on img2caption images for five epochs and save the best checkpoint. Finally, we get the well-trained Mol-VQGAN to tokenize molecule images.The original VQGAN will result in unclear and distorted images when encoding and decoding molecule images. After training, Mol-VQGAN can encode and decode molecule image almost the same with original image. Several examples are shown in Figure 17. The validation curve from start to the best checkpoint is shown in Figure 16.
@@
Figure 17: Examples of origin VQGAN and Mol-VQGAN. The origin images are shown in the first row. As shown in the second row, the origin VQGAN can not encode and decode molecule image clearly and accurately, resulting in distorted atoms and bonds that are hard to distinguish. As shown in the third row, our well-trained Mol-VQGAN can encode and decode molecule images with clear atoms and bonds, which is almost the same as the origin molecule images.
@@
ChemMLLM Supervised FineTuning Training For ChemMLLM training code, we utilize Lumina-mGPT framework (Liu et al. 2024). We train four variants ChemMLLM-7B, ChemMLLM-7B-pro2img, ChemMLLM-7B-img2img and ChemMLLM-34B. For ChemMLLM-7B, ChemMLLM-7B-pro2img and ChemMLLM-7B-img2img, we adopt Chameleon-7B as the base model. For ChemMLLM-34B, we adopt Lumina-mGPT-34B-512 as the base model.
@@
For ChemMLLM-7B, we train it on img2caption, img2property, img2smiles and property2img datasets for three epochs; For property2img task, we train and evaluate ChemMLLM-7B-pro2img. We find that directly training the model on raw data can not achieve good performance, so we do data augmentation by rotating images by $90^{\circ}$,$180^{\circ}$, and $270^{\circ}$ so as to generate augmented data that is four times the size of the original dataset. We train ChemMLLM-7B-pro2img on the augmented dataset for two epochs; For img2img task, we train and evaluate ChemMLLM-7B-img2img. we train it on img2img dataset for two epochs. For ChemMLLM-34B, we train it on all 5 tasks for two epochs and we report the best checkpoint result for each task (epoch 1.71 for the generation task, and epoch 2 for the understanding task). The training hyper-parameters primarily follow Lumina-mGPT (Liu et al. 2024).
@@
Model task # Training
ChemMLLM-7B img2caption 69,799
img2property 95,000
img2smiles 95,000
property2img 95,000
ChemMLLM-7B-pro2img property2img 380,000
ChemMLLM-7B-img2img img2img 157,673
ChemMLLM-34B img2caption 69,799
img2property 200,000
img2smiles 200,000
property2img 200,000
img2img 157,673
@@
Settings ChemMLLM ChemMLLM-pro2img ChemMLLM-img2img ChemMLLM-34B
z-loss weight 1e-05 1e-05 1e-05 1e-05
warmup epochs 0.01 0.01 0.01 0.01
learning rate 2e-05 2e-05 2e-05 2e-05
weight decay 0.1 0.1 0.1 0.1
drop rate 0.05 0.05 0.05 0.05
total bacth size 16×8×116$\times$ 8$\times$ 116 × 8 × 1 16×8×116$\times$ 8$\times$ 116 × 8 × 1 8×4×18$\times$ 4$\times$ 18 × 4 × 1 4×48×44$\times$ 48$\times$ 44 × 48 × 4
GPUs for training 8×8$\times$8 ×A800 (80G) 8×8$\times$8 × A800 (80G) 4×4$\times$4 × A800 (80G) 48×48$\times$48 × A800 (80G)
GPUs hours(h) 65 31 25 18,432